A Complete Bibliography of Publications in SIGCSE Bulletin: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

21 October 2023 Version 1.12

Title word cross-reference

x ≥ 2 [Bec20].

17th [Gri22].

2019 [FI20, Mur20, RG20]. 2020 [Ko20, MRMK20b, MRKM20, ZS20]. 2021 [Gar21, KMVH21, PS22, SB21a, SB21b, SMC^{+}20, SM21a, SMC^{+}21]. 2022 [BQ21, BQL22, BQ22, HVFF22, HA21, MD22b, MD22a, PJ22a, PJ22c, Set22b, Set22c, SDB22]. 2023 [DS22, DSD^{+}23a, DSD^{+}23b, FFDH23, LM23, LRCCP23, MJ23b, MJ23c, SSLM23]. 2024 [AA23, BRS23, Bec23a, SS23]. 202x [Bec20].

ACM [Bec23b, Bec23a, BC23, DL20, Dec20a, Dec21, Ste22]. across [PQ21]. Active [EL23]. address [MDM^{+}20, MDWP20, MDPR21, MW21, MWK22, MWWK22]. addressing [SIG20]. Adrienne [MWD22b]. After [BC23]. Alison [Dav20b]. American [MHD23]. Announcement [Bet23, McG22a]. Announcing [Bec23b, Bec23a]. Articulate [Ste22]. Artificial [Tou23]. Aspiring [AA23, HA21]. Attendees [DSD^{+}23a]. Award [Bec23a, Gar21, SD20, Bec23b]. awards [Mur20].

Binary [SWSM23]. Blazing [SS23]. Board [Cle23, CBD^{+}23, Set21, Set22a, Set22b, Set22c]. Boots [BQ22]. botch [McG22b]. bots [McG22b]. BPCnet.org [TWL20]. Briana [Mil20b]. Broadening [Bec23b, TWL20]. Bulletin [MW22].

Call [BRS23, DS22, LRCCP23, MJ23a, PJ22b, SSLM23, BQL22, FS20, Gri22, HVFF22, HA21, MD22a, OP21, PJ22a, PJ22c, SF21, SMC^{+}20, SDB22]. Calling [FI20, FS20, MJ23a, MJ23b, MJ23c, PJ22b, PJ22a, PJ22c, PS22, PJ23, SF21, OP21]. calls [DL20]. can [Bec20, MDM^{+}20, MDWP20, MDPR21, MW21, MWK22, MWWK22]. Celebrating [CGEM20]. Clear [Dav20b]. co [MW22]. co-editor [MW22]. Coding [MHD23]. committee [SIG20]. Common [PQ21]. CompEd [Bec20, BC23, LRCCP23]. Computational [Tou23]. Computing [Bec23b, Gri22, SWSM23, SIG20, TWL20]. Conference [PJ23, FI20, Ko20, MRMK20a, SF21]. Conferences [CGEM20]. Construct [CGEM20]. courses [DL20]. create [SIG20]. CS [McG22c, MDM^{+}20, MDWP20, MDPR21, MW21, MWK22, MWWK22, SM20, SM21b, SS23]. CS-Ed [SM20, SM21b]. CSEdGrad [Smi20]. CSEdResearch.org [McG22a, McG22b, McG22c]. csteachingtips.org [Lew20]. curricula [DL20].

Data [DL20]. Davis [Ano21b]. Dec [Bec20]. Decker [MWD22b]. DEIA [Cle23, CBD^{+}23]. Delay [BC23]. Developing [MHD23]. Development [AA23, HA21]. distinguished [Dec20a, Dec21]. Diverse [Ste22]. diversity [SIG20]. do [MDM^{+}20, MDWP20, MDPR21, MW21, MWK22, MWWK22]. don’t [McG22b].

earliest [Bec20]. Ed [SM20, SM21b]. editor [MW22]. Education [Bec23b, EL23, Gri22, SWSM23, SS23, SIG20]. educator [MDM^{+}20, MDWP20, MDPR21, MW21, MWK22, MWWK22]. Educators [AA23, HA21]. election [Set21, Set22b, Set22c]. Elementary [Tou23]. Employees [Ste22]. Equitable [Ste22]. Equity [SW23a, SIG20]. Everyone [CGEM20]. example [DL20]. executive [Set22a]. exist [MDM^{+}20, MDWP20, MDPR21, MW21, MWK22, MWWK22].

F2F [Bec20]. Fáilte [BQL22]. Fairness [EL23]. feedback [DL20]. Fellows [Dec21]. field [MDM^{+}20, MDWP20, MDPR21, MW21, MWK22, MWWK22]. Force [DL20]. Future [CGEM20, SGZ20]. game [Lew20]. gather [Bec20]. go [BQL22]. graduate [Smi20]. grants [Mur20]. Ground [BQ22]. groups [RG20].

Hearing [MHD23]. hÉirinn [BQL22].

ICER [FFDH23, HVFF22, Ko20, KMVH21, MRMK20a, MRMK20b, MRKM20]. Impaired [MHD23]. inclusion [SIG20]. Inclusive [Ste22]. inequity [MDM^{+}20, MDWP20, MDPR21, MW21, MWK22, MWWK22]. Information [DSD^{+}23a]. Instruction [MHD23]. Intelligence [Tou23]. Intelligent [Tou23]. International [SF21]. Introduce [Tou23]. Introducing [Set22a]. Invitation [MJ23b, MJ23c]. Involving [CGEM20]. issues [MDM^{+}20, MDWP20, MDPR21, MW21, MWK22, MWWK22]. ITiCSE [BQ21, BQL22, BQ22, CGEM20, LM23, RG20, SB21a, SB21b, SGZ20, SSLM23].

K12 [MHD23]. Karen [Ano21b]. Koli [FI20, FS20, MJ23a, MJ23b, MJ23c, OP21, PJ22b, PJ22a, PJ22c, PS22, PJ23, SF21].

Language [MHD23]. Leaders [Ste22]. Learning [EL23]. Leo [Dav21b]. let [McG22b]. Listening [Ste22]. Look [CGEM20].

Machine [EL23]. Member [Ano21a, Ano21b, Dav20a, Dav20b, Dav21a, Dav21b, Mil20a, Mil20b, MWD22b, MWV22, MWD22a, SWM22, SW23c, SWP23, SW23b, SWF23]. members [Dec20a, Dec21]. membership [PQ21]. microaggressions [Lew20]. Morrison [Mil20b].

named [Dec20a]. Need [Ste22]. Non [SWSM23]. Non-Binary [SWSM23]. November [Mur20].

online [OP21]. Organization [Ste22]. organizational [Dec20b].

Papers [LRCCP23, MJ23a, SSLM23, Gri22, SF21, SMC^{+}20]. Participation [Bec23b, DS22, PJ22b, BQL22, FS20, HVFF22, HA21, MD22a, OP21, PJ22a, PJ22c, TWL20]. Past [CGEM20, SGZ20]. Pearson [SW23c]. Podcast [Bet23, SM20, SM21b]. portal [TWL20]. Porter [Dav21b]. practice [McG22c]. Practices [EL23]. Preview [LM23, BQ21, KMVH21, MRMK20a, MRMK20b, SB21a, ZmS20]. Primary [Gri22]. problems [McG22c]. Professional [AA23, HA21]. Project [SWSM23]. projects [Mur20]. Promoting [EL23].

Recap [BQ22, FFDH23, SB21b, SMC^{+}21, ZS20]. Recipes [Bet23]. Reflective [EL23]. Report [DSD^{+}23b, PJ23, FI20, MD22b, PS22, RG20].

Research [SWSM23]. Resistance [Bet23]. resource [TWL20]. Resumes [BC23]. retrospective [MRKM20]. Robot [Tou23]. round [Mur20].

safely [Bec20]. School [Tou23]. Science [DL20]. season [SM21b]. Secondary [Gri22]. seeks [MW22]. serving [Smi20]. SIGCSE [Bec23b, Bec23a, AA23, BRS23, Cle23, CBD^{+}23, Dec20a, Dec20b, Dec21, DS22, DSD^{+}23a, DSD^{+}23b, Gar21, HA21, MD22b, MD22a, Mur20, SD20, Set21, Set22a, Set22b, Set22c, SMC^{+}20, SM21a, SMC^{+}21, SDB22, SS23, ZS20, ZmS20, MW22]. Sign [MHD23]. SIGs [PQ21]. special [Mur20]. Spotlight [SWM22, SW23a, SW23c, SWP23, SW23b, SWF23, Ano21a, Ano21b, Dav20a, Dav20b, Dav21a, Dav21b, Mil20a, Mil20b, MWD22b, MWV22, MWD22a]. Students [MHD23, Tou23, Smi20]. Submissions [BRS23, PJ22c, SDB22]. support [TWL20]. supporting [Smi20]. surveys [McG22b]. Symposium [BRS23, DS22, DSD^{+}23a, DSD^{+}23b, MD22a, SMC^{+}20, SWEH22, SDB22, SS23, SD20, SM21a, SMC^{+}21, ZmS20].

Tamara [SW23c]. Task [DL20]. teacher [McG22c]. teaching [McG22c]. Technical [BRS23, DS22, DSD^{+}23a, DSD^{+}23b, MD22a, SMC^{+}20, SWEH22, SDB22, SS23, SD20, SM21a, SMC^{+}21, ZmS20]. think [MDM^{+}20, MDWP20, MDPR21, MW21, MWK22, MWWK22]. Thinking [Tou23]. Trails [SS23]. Trans [SWSM23]. travel [SD20]. TS [AA23, MD22b]. Two [BC23]. Two-year [BC23].

upcoming [McG22c]. Update [Cle23, CBD^{+}23]. updates [Dec20b]. Using [MHD23, Tou23].

Videos [MHD23]. virtual [Ko20].

webinar [McG22c]. Winners [Bec23a, Gar21, SD20]. WiPSCE [Gri22]. Within [Ste22]. work [TWL20]. working [RG20]. Workshop [AA23, Gri22, HA21].

year [Bec20, BC23].

References

Allen:2023:PDW

[AA23] Meghan Allen and Mohammad Azhar. Professional development workshop for new and aspiring educators at SIGCSE TS 2024. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(4):3–4, October 2023. CODEN SIGSD3. ISSN

           0097-8418. URL https://dl.acm.org/doi/10.1145/3629712.
           3629715.

Anonymous:2021:MS

[Ano21a] Anonymous. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(4):6–9, October 2021. CODEN SIGSD3. ISSN 0097-8418. URL https:/

/dl.acm.org/doi/10.1145/3505166.3505169.

Anonymous:2021:MSK

[Ano21b] Anonymous. Member spotlight: Karen Davis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(3):6–9, July 2021. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3483403.3483409.

Becker:2023:ACR

[BC23] Brett A. Becker and Venkatesh Choppella. ACM CompEd resumes after two-year delay! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(1):6, January 2023. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/

doi/10.1145/3584667.3584673.

Becker:2020:CDX

[Bec20] Brett Becker. CompEd: Dec 202x (x ≥ 2) … not 2021, but the earliest year we can gather safely F2F). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(4):4, October 2020. CODEN SIGSD3. ISSN 0097-8418. URL https:/

/dl.acm.org/doi/10.1145/3430905.3430907.

Becker:2023:AASb

[Bec23a] Brett Becker. Announcing the ACM SIGCSE 2024 award winners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(4):2, October 2023. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3629712.
3629713.

Becker:2023:AASa

[Bec23b] Brett Becker. Announcing the ACM SIGCSE Broadening Participation in Computing Education Award. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(3): 2–3, July 2023. CODEN SIGSD3. ISSN 0097-8418. URL https:

//dl.acm.org/doi/10.1145/3610585.3610586.

Bettin:2023:RRP

[Bet23] Briana Bettin. Recipes for resistance podcast announcement. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(1):4–5, January 2023. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3584667.
3584671.

Becker:2021:IP

[BQ21] Brett Becker and Keith Quille. ITiCSE 2022 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(4):3–6, October 2021. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3505166.
3505168.

Becker:2022:IRB

[BQ22] Brett Becker and Keith Quille. ITiCSE 2022 recap: Boots on the ground! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(4):2–3, October 2022. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3577871.3577872.

Becker:2022:ICP

[BQL22] Brett Becker, Keith Quille, and Mikko-Jussi Laakso. ITiCSE 2022 call for participation: Fáilte go hÉirinn! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(2): 5–6, April 2022. CODEN SIGSD3. ISSN 0097-8418. URL https:

//dl.acm.org/doi/10.1145/3538522.3538526.

Battestilli:2023:STS

[BRS23] Lina Battestilli, Samuel Rebelski, and Libby Shoop. SIGCSE Technical Symposium 2024: Call for submissions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(3):3, July 2023. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3610585.3610587.

Clear:2023:SBDb

[CBD^{+}23] Alison Clear, Brett A. Becker, Jill Denner, Dan Garcia, Rodrigo Duran, Yolanda A. Rankin, Judy Sheard, and Adrienne Decker. SIGCSE Board DEIA update. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(2):2–3, April 2023. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.

org/doi/10.1145/3607479.3607480.

Cassel:2020:CIC

[CGEM20] Lillian (Boots) Cassel, Judith Gal-Ezer, and Francesco Maiorana. Celebrating 25 ITiCSE conferences: Involving everyone to look at the past to construct the future. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(3):6–7, August 2020. CODEN SIGSD3. ISSN 0097-8418. URL https:/

/dl.acm.org/doi/10.1145/3419040.3419045.

Clear:2023:SBDa

[Cle23] Alison Clear. SIGCSE Board DEIA update. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(1):2, January 2023. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3584667.3584668.

Davis:2020:MS

[Dav20a] Karen Davis. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(3):10, August 2020. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.

org/doi/10.1145/3419040.3419047.

Davis:2020:MSA

[Dav20b] Karen Davis. Member spotlight: Alison Clear. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(1):13–14, January 2020. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3380469.3380478.

Davis:2021:MS

[Dav21a] Karen Davis. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(1):7–9, January 2021. CODEN SIGSD3. ISSN 0097-8418. URL https:/

/dl.acm.org/doi/10.1145/3457573.3457579.

Davis:2021:MSL

[Dav21b] Karen Davis. Member spotlight: Leo Porter. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(2):7–10, April 2021. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3466995.3467000.

Decker:2020:SMN

[Dec20a] Adrienne Decker. SIGCSE members named distinguished ACM members. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(1):2, January 2020. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3380469.3380470.

Decker:2020:SOU

[Dec20b] Adrienne Decker. SIGCSE organizational updates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(2):2–4, April 2020. CODEN SIGSD3. ISSN 0097-

8418. URL https://dl.acm.org/doi/abs/10.1145/3397568.
3397569.

Decker:2021:AFD

[Dec21] Adrienne Decker. ACM Fellows and distinguished members in SIGCSE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(1):2–3, January 2021. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3457573.3457574.

Danyluk:2020:ADS

[DL20] Andrea Danyluk and Paul Leidig. ACM Data Science Task Force calls for curricula feedback and example courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(2):8, April 2020. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/abs/10.1145/3397568.3397571.

Doyle:2022:STS

[DS22] Maureen Doyle and Ben Stephenson. SIGCSE Technical Symposium 2023: Call for participation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(4):4–5, October 2022. CODEN SIGSD3. ISSN 0097-8418. URL https:/

/dl.acm.org/doi/10.1145/3577871.3577874.

Doyle:2023:STSa

[DSD^{+}23a] Maureen Doyle, Ben Stephenson, Brian Dorn, Leen-Kiat Soh, Lina Battestilli, Kristin Stephens-Martinez, and Delaram Yazdansepas. SIGCSE Technical Symposium 2023: Information for attendees. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(1):2–4, January 2023. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3584667.
3584669.

Doyle:2023:STSb

[DSD^{+}23b] Maureen Doyle, Ben Stephenson, Brian Dorn, Leen-Kiat Soh, Lina Battestilli, Kristin Stephens-Martinez, and Delaram Yazdansepas. SIGCSE Technical Symposium 2023 report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(2):3–6, April 2023. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3607479.3607481.

Elglaly:2023:PML

[EL23] Yasmine N. Elglaly and Yudong Liu. Promoting machine learning fairness education through active learning and reflective practices. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(3):4–6, July 2023. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3610585.
3610589.

Franklin:2023:IR

[FFDH23] Diana Franklin, Kathi Fisler, Paul Denny, and Margaret Hamilton. ICER 2023 recap. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(4):5, October 2023. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/

doi/10.1145/3629712.3629717.

Falkner:2020:KCCa

[FI20] Nickolas Falkner and Petri Ihantola. Koli Calling 2019 conference report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(1):11–12, January 2020. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3380469.3380476.

Falkner:2020:KCCb

[FS20] Nick Falkner and Otto Seppälä. Koli calling: 2020 call for participation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(3):4–5, August 2020. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3419040.3419043.

Garcia:2021:SAW

[Gar21] Dan Garcia. SIGCSE 2021 award winners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(1):4, January 2021. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3457573.3457575.

Grillenberger:2022:WPS

[Gri22] Maureen Grillenberger. 17th Workshop in Primary and Secondary Computing Education (WiPSCE 2022) call for papers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(1):6–7, January 2022. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3516410.
3516415.

Horton:2021:PDW

[HA21] Diane Horton and Meghan Allen. Professional development workshop for new and aspiring educators at SIGCSE 2022: call for participation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(4):3, October 2021. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3505166.3505167.

Hauswirth:2022:ICP

[HVFF22] Matthias Hauswirth, Jan Vahrenhold, Diana Franklin, and Kathi Fisler. ICER 2022 call for participation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(2):6, April 2022. CODEN SIGSD3. ISSN 0097-8418. URL https://

dl.acm.org/doi/10.1145/3538522.3538527.

Ko:2021:IP

[KMVH21] Amy Ko, Renée McCauley, Jan Vahrenhold, and Matthias Hauswirth. ICER 2021 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(1):6, January 2021. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.

org/doi/10.1145/3457573.3457577.

Ko:2020:IVC

[Ko20] Amy J. Ko. ICER 2020 virtual conference. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(3):3–4, August 2020. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3419040.3419042.

Lewis:2020:NCO

[Lew20] Colleen Lewis. New from csteachingtips.org: microaggressions: the game! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(1):10, January 2020. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3380469.3380475.

Laakso:2023:IP

[LM23] Mikko-Jussi Laakso and Mattia Monga. ITiCSE 2023 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(2):6–7, April 2023. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3607479.
3607482.

Luxton-Reilly:2023:CCP

[LRCCP23] Andrew Luxton-Reilly, Michelle Craig, Venkatesh Choppella, and BD Phatak. CompEd 2023 call for papers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(2):7–8, April 2023. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3607479.3607483.

McGill:2022:COA

[McG22a] Monica McGill. CSEdResearch.org announcement. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(1):6, January 2022. CODEN SIGSD3. ISSN 0097-8418.

URL https://dl.acm.org/doi/10.1145/3516410.3516414.

McGill:2022:COD

[McG22b] Monica McGill. CSEdResearch.org: don’t let bots botch your surveys. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(2):8, April 2022. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3538522.
3538529.

McGill:2022:COU

[McG22c] Monica McGill. CSEdResearch.org: upcoming webinar on teachers’ problems of practice teaching CS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(3):5, July 2022. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.

acm.org/doi/10.1145/3551627.3551631.

Merkle:2022:STS

[MD22a] Larry Merkle and Maureen Doyle. SIGCSE Technical Symposium 2022 call for participation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(1):4–5, January 2022. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.

org/doi/10.1145/3516410.3516412.

Merkle:2022:STR

[MD22b] Larry Merkle and Maureen Doyle. SIGCSE TS 2022 report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(2):3–5, April 2022. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3538522.
3538525.

Miller:2020:CEHa

[MDM^{+}20] Jeffrey Miller, Karen C. Davis, Brandeis Marshall, Nicki Washington, Manuel A. Pérez-Quiñones, Leo Porter, and Juan E. Gilbert. As a CS educator, how do you think we can address inequity issues that exist in the field? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(3):7–10, August 2020. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/

doi/10.1145/3419040.3419046.

Miller:2021:CEHa

[MDPR21] Jeffrey Miller, Karen C. Davis, Fay Cobb Payton, and Saty Raghavachary. As a CS educator, how do you think we can address inequity issues that exist in the field? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(2):6–7, April 2021. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3466995.3466999.

Miller:2020:CEHb

[MDWP20] Jeffrey Miller, Karen C. Davis, Shana V. White, and Robert Parke. As a CS educator, how do you think we can address inequity issues that exist in the field? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(4):8–9, October 2020. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/

doi/10.1145/3430905.3430911.

Marghitu:2023:DCI

[MHD23] Daniela Marghitu, Ayanna Howard, and Meenakshi Das. Developing coding instruction videos for K12 hearing impaired students using American Sign Language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(2):10–12, April 2023. CODEN SIGSD3. ISSN 0097-8418. URL https://

dl.acm.org/doi/10.1145/3607479.3607486.

Miller:2020:MS

[Mil20a] Jeffrey Miller. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(4):9–11, October 2020. CODEN SIGSD3. ISSN 0097-8418. URL https:/

/dl.acm.org/doi/10.1145/3430905.3430912.

Miller:2020:MSB

[Mil20b] Jeffrey Miller. Member spotlight: Briana Morrison. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(2):13–14, April 2020. CODEN SIGSD3. ISSN 0097-

8418. URL https://dl.acm.org/doi/abs/10.1145/3397568.
3397574.

Muhling:2023:KCC

[MJ23a] Andreas Mühling and Ilkka Jormanainen. Koli calling 2023 call for papers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(2):8, April 2023. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3607479.3607484.

Muhling:2023:KCIa

[MJ23b] Andreas Mühling and Ilkka Jormanainen. Koli Calling 2023 invitation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(3):3–4, July 2023. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3610585.
3610588.

Muhling:2023:KCIb

[MJ23c] Andreas Mühling and Ilkka Jormanainen. Koli Calling 2023 invitation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(4):4, October 2023. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3629712.3629716.

Moskal:2020:IR

[MRKM20] Adon Moskal, Anthony Robins, Amy Ko, and Renee McCauley. ICER 2020 retrospective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(4):7–8, October 2020. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/

doi/10.1145/3430905.3430910.

Moskal:2020:ICP

[MRMK20a] Adon Moskal, Anothony Robins, Renee McCauley, and Amy J. Ko. ICER 2020 conference preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(1):9–10, January 2020. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-

3927 (electronic). URL https://dl.acm.org/doi/abs/10.1145/
3380469.3380474.

Moskal:2020:IP

[MRMK20b] Adon Moskal, Anothony Robins, Renee McCauley, and Amy J. Ko. ICER 2020 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(2):10, April 2020.

CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/

doi/abs/10.1145/3397568.3397573.

Murphy:2020:SSP

[Mur20] Laurie Murphy. SIGCSE special projects grants awards: November 2019 round. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(2):9–10, April 2020. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/abs/
10.1145/3397568.3397572.

Miller:2021:CEHb

[MW21] Jeffrey Miller and Charles Wallace. As a CS educator, how do you think we can address inequity issues that exist in the field? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(3):6, July 2021. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3483403.
3483408.

Miller:2022:SBS

[MW22] Jeffrey Miller and Charles Wallace. SIGCSE Bulletin seeks a new co-editor. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(2):3, April 2022. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3538522.3538524.

Miller:2022:MSc

[MWD22a] Jeffrey Miller, Charles Wallace, and Mats Daniels. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(3):6–10, July 2022. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3551627.3551633.

Miller:2022:MSA

[MWD22b] Jeffrey Miller, Charles Wallace, and Adrienne Decker. Member spotlight: Adrienne Decker. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(2):8–11, April 2022. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.

org/doi/10.1145/3538522.3538530.

Miller:2022:CEHa

[MWK22] Jeffrey Miller, Charles Wallace, and Maria Klawe. As a CS educator, how do you think we can address inequity issues that exist in the field? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(1):7, January 2022. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3516410.3516416.

Miller:2022:MSb

[MWV22] Jeffrey Miller, Charles Wallace, and Frank Vahid. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(1):8–9, January 2022. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3516410.3516417.

Miller:2022:CEHb

[MWWK22] Jeffrey Miller, Charles Wallace, Kendra Walther, and Shalini Kesar. As a CS educator, how do you think we can address inequity issues that exist in the field? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(3):5, July 2022. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/

doi/10.1145/3551627.3551632.

Otto:2021:KCO

[OP21] Seppälä Otto and Andrew Petersen. Koli Calling is online: call for participation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(3):3, July 2021. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3483403.3483404.

Petersen:2022:KCCa

[PJ22a] Andrew Petersen and Ilkka Jormanainen. Koli Calling 2022 call for participation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(2):6–8, April 2022. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3538522.3538528.

Petersen:2022:KCC

[PJ22b] Andrew Petersen and Ilkka Jormanainen. Koli calling 2022: Call for participation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(4):3–4, October 2022. CO-DEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/

10.1145/3577871.3577873.

Petersen:2022:KCCb

[PJ22c] Andrew Petersen and Ilkka Jormanainen. Koli Calling 2022: call for submissions and participation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(3):4, July 2022. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.

org/doi/10.1145/3551627.3551629.

Petersen:2023:KCC

[PJ23] Andrew Petersen and Ilkka Jormanainen. Koli calling conference report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(1):5–6, January 2023. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3584667.3584672.

Perez-Quinones:2021:CMA

[PQ21] Manuel Pérez-Quiñones. Common membership across SIGs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(2):3, April 2021. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3466995.
3466996.

Petersen:2022:KCR

[PS22] Andrew Petersen and Otto Seppälä. Koli Calling 2021 report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(1):3–4, January 2022. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3516410.
3516411.

Rossling:2020:IWG

[RG20] Guido Rößling and Michail Giannakos. ITiCSE 2019 working groups report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(1):2–4, January 2020. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3380469.3380471.

Schulte:2021:IP

[SB21a] Carsten Schulte and Brett Becker. ITiCSE 2021 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(2):5, April 2021. CODEN SIGSD3. ISSN 0097-8418.

URL https://dl.acm.org/doi/10.1145/3466995.3466998.

Schulte:2021:IR

[SB21b] Carsten Schulte and Brett Becker. ITiCSE 2021 recap. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(3):3–4, July 2021. CODEN SIGSD3. ISSN 0097-8418.

URL https://dl.acm.org/doi/10.1145/3483403.3483405.

Settle:2020:STS

[SD20] Amber Settle and Karen Davis. SIGCSE technical symposium 2020 travel award winners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(1):6–8, January 2020. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3380469.
3380473.

Soh:2022:STS

[SDB22] Leen-Kiat Soh, Brian Dorn, and Lina Battestilli. SIGCSE Technical Symposium 2022: call for submissions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(3):4–5, July 2022. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3551627.3551630.

Settle:2021:SBE

[Set21] Amber Settle. 2022 SIGCSE board election. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(3):5, July 2021. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3483403.3483407.

Settle:2022:ISE

[Set22a] Amber Settle. Introducing the 2022–2025 SIGCSE executive board. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(3):3–4, July 2022. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3551627.3551628.

Settle:2022:SBEa

[Set22b] Amber Settle. SIGCSE 2022 board election. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(1):5, January 2022. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3516410.3516413.

Settle:2022:SBEb

[Set22c] Amber Settle. SIGCSE 2022 board election. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(2):2–3, April 2022. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3538522.3538523.

Seppala:2021:KCI

[SF21] Otto Seppälä and Nickolas Falkner. Koli Calling International Conference call for papers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(2):3–4, April 2021. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.

org/doi/10.1145/3466995.3466997.

Settle:2020:PFI

[SGZ20] Amber Settle, Michail Giannakos, and Mark Zarb. The past and future of ITiCSE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(4):3–4, October 2020. CO-DEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/

10.1145/3430905.3430906.

Board:2020:SBC

[SIG20] SIGCSE Board (2019–2022). SIGCSE board to create committee addressing equity, diversity, and inclusion in computing education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(3):3, August 2020. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3419040.
3419041.

Stephens-Martinez:2020:CEP

[SM20] Kristin Stephens-Martinez. The CS-Ed podcast. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52 (1):12, January 2020. CODEN SIGSD3. ISSN 0097-8418 (print),

2331-3927 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3380469.3380477.

Sherriff:2021:STSa

[SM21a] Mark Sherriff and Larry Merkle. SIGCSE 2021 technical symposium. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(1):5, January 2021. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3457573.3457576.

Stephens-Martinez:2021:CEP

[SM21b] Kristin Stephens-Martinez. The CS-Ed podcast season 2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(1):6–7, January 2021. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3457573.
3457578.

Sherriff:2020:STS

[SMC^{+}20] Mark Sherriff, Larry Merkle, Pam Cutter, Alvaro Monge, and Judy Sheard. SIGCSE Technical Symposium 2021 call for papers.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(3):5–6, August 2020. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3419040.
3419044.

Sherriff:2021:STSb

[SMC^{+}21] Mark Sherriff, Larry Merkle, Pam Cutter, Alvaro Monge, and Judy Sheard. SIGCSE 2021 technical symposium recap. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 53(3):4–5, July 2021. CODEN SIGSD3. ISSN 0097-8418.

URL https://dl.acm.org/doi/10.1145/3483403.3483406.

Smith:2020:CSS

[Smi20] Julie Smith. CSEdGrad: @serving and supporting graduate students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(4):4–6, October 2020. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3430905.3430908.

Stephenson:2023:STS

[SS23] Ben Stephenson and Jeffrey Stone. SIGCSE Technical Symposium 2024: Blazing new trails in CS education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(4): 2–3, October 2023. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3629712.3629714.

Simon:2023:ICP

[SSLM23] Simon, Judy Sheard, Mikko-Jussi Laakso, and Mattia Monga. ITiCSE 2023 call for papers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(1):4, January 2023. CODEN SIGSD3. ISSN 0097-8418. URL https://dl.acm.

org/doi/10.1145/3584667.3584670.

Stephenson:2022:LWA

[Ste22] Chris Stephenson. Listening within: ACM employees and leaders articulate the need for a more diverse, equitable, and inclusive organization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(4):5–7, October 2022. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3577871.3577875.

Smith:2023:ES

[SW23a] Julie M. Smith and Charles Wallace. Equity spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(2):12–13, April 2023. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3607479.
3607487.

Smith:2023:MSb

[SW23b] Julie M. Smith and Charles Wallace. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(2):13–16, April 2023. CODEN SIGSD3. ISSN

0097-8418. URL https://dl.acm.org/doi/10.1145/3607479.
3607488.

Smith:2023:MST

[SW23c] Julie M. Smith and Charles Wallace. Member spotlight: Tamara Pearson. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(4):6–7, October 2023. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3629712.3629718.

Smith:2022:TSA

[SWEH22] Julie M. Smith, Charles Wallace, Yasmine Elglaly, and Earl W. Huff. A technical symposium for all. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(4): 7–9, October 2022. CODEN SIGSD3. ISSN 0097-8418. URL

https://dl.acm.org/doi/10.1145/3577871.3577876.

Smith:2023:MSc

[SWF23] Julie M. Smith, Charles Wallace, and Casey Fiesler. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(3):13–16, July 2023. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3610585.3610590.

Smith:2022:MS

[SWM22] Julie M. Smith, Charles Wallace, and Jeffrey Miller. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 54(4):9–11, October 2022. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3577871.3577877.

Smith:2023:MSa

[SWP23] Julie M. Smith, Charles Wallace, and Anne-Kathrin Peters. Member spotlight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(1):8–12, January 2023. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3584667.3584675.

Smith:2023:TNB

[SWSM23] Julie M. Smith, Charles Wallace, Stacey Sexton, and Amanda Menier. The Trans & Non-Binary Computing Education Research Project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(1):6–8, January 2023. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3584667.3584674.

Touretzky:2023:UIR

[Tou23] David S. Touretzky. Using an intelligent robot to introduce elementary school students to computational thinking and artificial intelligence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 55(2):8–10, April 2023. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3607479.3607485.

Tamer:2020:BOR

[TWL20] Burçin Tamer, Heather Wright, and Colleen Lewis. BPCnet.org: a resource portal to support your broadening participation in computing work. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(4):3–4, October 2020. CODEN

SIGSD3. ISSN 0097-8418. URL https://dl.acm.org/doi/10.
1145/3430905.3430909.

Zhang:2020:STS

[ZmS20] Jian Zhang and mark Sherriff. SIGCSE technical symposium 2020 preview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(1):5–6, January 2020. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3380469.3380472.

Zhang:2020:SR

[ZS20] Jian Zhang and Mark Sherriff. SIGCSE 2020 recap. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 52(2):4–7, April 2020. CODEN SIGSD3. ISSN 0097-

8418. URL https://dl.acm.org/doi/abs/10.1145/3397568.
3397570.